OpenStudy (anonymous):

*graph the function* F(x) = {x + 3, if x < 0 { 2x, if x > 0

OpenStudy (nincompoop):

let us see you try to figure it out first; this way so that we know what part you are struggling about

OpenStudy (anonymous):



OpenStudy (anonymous):

The way my teacher explained it was like X<0 X>0 Y= x+3 Y=2x M= 1/1 M=2/1 =2 B=3 B= 0 But I'm still like a little confused

OpenStudy (anonymous):

what is B?

OpenStudy (anonymous):

I'm thinking is he using the y=mx+b?...

OpenStudy (anonymous):

oh yes that's true.. so what makes you confused? am i wrong with that graph?

OpenStudy (anonymous):

It's just the whole figuring out how to set it up correctly. No the graph I think is right.

OpenStudy (nincompoop):